OpenStudy (anonymous):

A partial solution to the equation 8b + 12 = 5(4 + 2b) is given below. Choose the step that is missing from the solution. 8b + 12 = 5(4 + 2b) 8b = 8 + 10b −2b = 8 b = −4 8b + 12 = 20 + 2b 8b + 12 = 20 + 10b 8b + 12 = 9 + 7b 8b + 12 = 4 + 10b

OpenStudy (anonymous):

8b +12 = 20 + 10b

OpenStudy (anonymous):

8b + 12 = 5(4 + 2b) 8b+12=20+10b 8b-10b=20-12 -2b=8 b=8/-2 b=-4

Parth (parthkohli):

My answer is wrong. Sorry :(

OpenStudy (anonymous):

Thanks for trying lol